OpenStudy (anonymous):

need help Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0. A.)slope = −1/4 and y-intercept = −9/4 B.)slope = 1/4 and y-intercept = −9/4 C.)slope = 1/4 and y-intercept = 9/4 D.)slope = −1/4 and y-intercept = 9/4

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So take you equation and put it in y = mx + b form. Can you do that ?

OpenStudy (anonymous):

would it be D cause the y intercept is positive? or is the y intercept positive? I already got the slope if its right but don't know if the y intercept positive or negative

OpenStudy (texaschic101):

x - 4(y + 1) - 5 = 0 -- distribute through the parenthesis x - 4y - 4 - 5 = 0 -- simplify x - 4y - 9 = 0 -- add 9 to both sides x - 4y = 9 -- you can rearrange this if it makes it easier -4y + x = 9 -- now multiply by -1 to make x positive 4y - x = 9 --- add x to both sides 4y = x + 9 -- divide both sides by 4 y = 1/4x + 9/4 so the slope(m) = 1/4 and the y intercept = 9/4

OpenStudy (anonymous):

I must of mest up cause I thought the slope was negative

OpenStudy (texaschic101):

it can get confusing at times